# Building a Correct-by-Design Lakehouse

### v2

@@ Abstract.
Lakehouses are the default cloud platform for analytics and AI, but they become unsafe when untrusted actors concurrently operate on production data: upstream–downstream mismatches surface only at runtime, and multi-table pipelines can leak partial effects. Inspired by software engineering, we design Bauplan, a code-first lakehouse that aims to make (most) illegal states unrepresentable using familiar abstractions. Bauplan acts along three axes: typed table contracts to make pipeline boundaries checkable, Git-like data versioning for review and reproducibility, and transactional runs that guarantee pipeline-level atomicity. We report early results from a lightweight formal transaction model and discuss future work motivated by counterexamples.
@@ 3. Design principles
Accordingly, we first introduce typed contracts to make pipeline interfaces machine-checkable. We then show how mapping table snapshots to Git-like commits/branches enables review and reproducibility. Finally, we define run semantics that publish multi-table effects atomically via branch-and-merge. After our core design choices have been described, Section 4 will provide a deep dive on the correctness primitives.
@@ 3.3. Transactional pipelines
Creating MVCC-style (Bernstein et al., 1987) transaction boundaries in the lakehouse is not straightforward. Unlike vertically integrated databases, lakehouse pipelines run on ephemeral compute and decoupled storage. Bauplan DAGs are restricted to use functions with signature Table(s) -¿ Table, but within those functions, anything can happen: any SQL query or Python version and package are allowed, as long as the signature is respected. At the programming level, we still want to pursue our general principle: we don’t want actors to learn new complex abstractions, nor do we want to delegate correctness to the application layer. Our key insight is to slightly modify the semantics of the run API and logically couple function execution with data branches, as exemplified by the following protocol. If $B$ is the target branch for execution, a new run:
@@ 5. Related work
Our work aligns with recent database literature highlighting similarities between Git primitives and transactions. While TARDIS (Crooks et al., 2016) proposes branch-and-merge for weak isolation in a key-value store, Bauplan’s APIs apply to distributed storage in a multilanguage OLAP setting, re-using collaboration abstractions for transaction boundaries. While previous work suggests that Git commit graph is a form of transaction (Gunce Su Yilmaz, 2025), our contribution targets a different correctness boundary: data pipelines are multi-table and multi-language, and the dominant failure mode is pipeline-level partial publication rather than tuple-level anomalies. Moreover, our Alloy counterexample highlights that the additional expressivity enabled by nested branching can be a safety hazard for pipeline transactions unless transactional branches are given stricter visibility. As part of future work, we plan on modelling isolation levels in Bauplan following the formalization in (Cerone et al., 2015).

### v3

@@ Abstract.
Lakehouses are now the default substrate for analytics and AI, but they remain fragile under concurrent, untrusted change: schema mismatches often surface only at runtime, development and production easily diverge, and multi-table pipelines can expose partial results after failure. We present Bauplan, a code-first lakehouse that aims to eliminate a broad class of these failures by construction. Bauplan builds on a storage substrate that already provides atomic single-table snapshot evolution, and adds three pipeline-level correctness mechanisms: typed table contracts to make transformation boundaries checkable, Git-like data versioning to support reproducible collaboration and review, and transactional runs that guarantee atomic publication of an entire pipeline execution. We describe the system design, show how these abstractions fit together into a unified programming model for humans and agents, and report early results from a lightweight Alloy model that both validates key intuitions and exposes subtle counterexamples around transactional branch visibility. Our experience suggests that correctness in the lakehouse is best addressed not by patching failures after the fact, but by restricting the programming model so that many illegal states become unrepresentable.
@@ 3. Design principles
Importantly, Bauplan is layered on top of storage that already provides atomic single-table snapshot evolution and optimistic concurrency control (e.g., Apache Iceberg plus a catalog/metadata service). We do not re-prove those guarantees in this work, but assume they work as expected. Bauplan adds three higher-level correctness properties at the pipeline layer: checkable interfaces between transformations, reproducible multi-table collaboration through versioned branches and commits, and atomic publication of an entire pipeline run. In other words, Iceberg provides table-level atomicity; Bauplan provides pipeline-level structure and publication semantics. After our core design choices have been described, Section 4 will provide a deep dive on the correctness primitives.
@@ 3.3. Transactional pipelines
Creating MVCC-style (Bernstein et al., 1987) transaction boundaries in the lakehouse is not straightforward. Unlike vertically integrated databases, lakehouse pipelines run on ephemeral compute and decoupled storage. Bauplan DAGs are restricted to use functions with signature Table(s) -¿ Table (in line with other data engineering frameworks (dbt-labs, 2024)), but within those functions, anything can happen: any SQL query or Python version and package are allowed, as long as the signature is respected. At the programming level, we still want to pursue our general principle: we don’t want actors to learn new complex abstractions, nor do we want to delegate correctness to the application layer. Our key insight is to slightly modify the semantics of the run API and logically couple function execution with data branches, as exemplified by the following protocol. If $B$ is the target branch for execution, a new run:
@@
The transactional branch protocol introduces metadata and coordination overhead relative to direct writes to the target branch. In our target OLAP setting, this trade-off is acceptable because pipelines are coarse-grained, multi-table jobs whose dominant costs are typically storage I/O and compute, not branch creation or metadata updates. The protocol may be less attractive for extremely high-frequency or fine-grained update workloads; this is a deliberate design choice, and further distinguishes our target from OLTP branching systems.
@@ 5. Related work
Our work aligns with recent database literature highlighting similarities between Git primitives and transactions. While TARDIS (Crooks et al., 2016) proposes branch-and-merge for weak isolation in a key-value store, Bauplan’s APIs apply to distributed storage in a multilanguage OLAP setting, re-using collaboration abstractions for transaction boundaries. Branching systems for OLTP databases (DoltHub, 2026) are related but address a different problem, as they operate over transactional row-store workloads.
While previous work suggests that Git commit graph is a form of transaction (Gunce Su Yilmaz, 2025), our contribution targets a different correctness boundary: data pipelines are multi-table and multi-language, and the dominant failure mode is pipeline-level partial publication rather than tuple-level anomalies. Moreover, our Alloy counterexample highlights that the additional expressivity enabled by nested branching can be a safety hazard for pipeline transactions unless transactional branches are given stricter visibility. As part of future work, we plan on modelling isolation levels in Bauplan following the formalization in (Cerone et al., 2015).